OpenStudy (anonymous):

How many solutions (x,y,z) that satisfies xyz=1000 with x,y, and z are positive integer numbers ?

OpenStudy (shubhamsrg):

1000 = 2^3 * 5^3 = 2*2*2 *5*5*5 *1 you have to form 3 sets from these 7 elements such that all elements are used. question is how many such formations are possible such that none is empty set..

OpenStudy (shubhamsrg):

each set would represent x,y or z

OpenStudy (anonymous):

still confused --"